OpenStudy (loser66):

\[\int \dfrac{1}{x^2\sqrt{x^2-5}}dx\] Please, help

OpenStudy (loser66):

We have \(tan^2t = sec^2 t-1\) and \(cot^2t = csc^2 t -1\) My question: why can't we use the second one to solve this problem?

OpenStudy (loser66):

Thanks for the link @inkyvoyd . But when we use the second formula to solve it, we get a different answer. why?

OpenStudy (inkyvoyd):

What answer do you get?

OpenStudy (loser66):

identity sin^2 t + cos^2 t =1 also.

jhonyy9 (jhonyy9):

like a first step rewrite x^2sqrt(x^2 -5) = sqrt(x^6 -5x4) than note sqrt(x^6 -5x^4) = u

OpenStudy (loser66):

@jhonyy9 we go nowhere then.

OpenStudy (loser66):

Ok, I show my work for the first formula

OpenStudy (loser66):

Let \(x =\sqrt5 sec t\) then \(x^2 = 5sec^2 t\) and \(\sqrt{x^2-5}=\sqrt{5(sec^2t -1)}=tant \sqrt5\) and then \(dx= \sqrt5 sec t tant dt\) It becomes \[\int \dfrac{\sqrt5 sect t tant dt}{5sec^2t tant\sqrt5}=\dfrac{1}{5}\int cos tdt=\dfrac{1}{5}sin t+C\]

OpenStudy (inkyvoyd):

then x^2=25sec^2 t...

OpenStudy (holsteremission):

@Loser66 Nothing is wrong with using one formula over the other, but the second introduces a minus sign which could potentially lead to a mistake if you're not too careful.

OpenStudy (loser66):

@HolsterEmission I will show you the second formula

OpenStudy (loser66):

Please, help me to check out the mistake

OpenStudy (loser66):

The step to convert from sin t to x is easy. So that I don't waste your time on it. But the second formula leads me to cos x /5 That is my problem

OpenStudy (loser66):

Now, if let \(x =\sqrt 5 csc t\), then \(x^2 = 5 csc^2 t\) and \(\sqrt{x^2 -5}= cot t\sqrt 5\) \(dx=-\sqrt5 csc t cot t dt\) put them all \[\int \dfrac{-\sqrt5 csct cot t dt}{5\sqrt5csc^2t cott}=-\dfrac{1}{5}\int \dfrac{1}{csc t}dt=\dfrac{-1}{5}\int sin tdt=\dfrac{1}{5}cost +C\]

OpenStudy (loser66):

Where is my mistake?

OpenStudy (agent0smith):

You need to convert back to the original substitution... I'm willing to bet you'll see they're equivalent once you do that.

OpenStudy (holsteremission):

No mistake. With the second substitution, the last step would be to find \(\cos\left(\mathrm{arccsc}\frac{x}{\sqrt5}\right)\), while the first substitution had you find \(\sin\left(\mathrm{arcsec}\frac{x}{\sqrt5}\right)\).

OpenStudy (loser66):

@agent0smith what is the triangle you use to convert?

OpenStudy (agent0smith):

"The step to convert from sin t to x is easy. So that I don't waste your time on it" This is your mistake. Assuming that they're wrong because you didn't complete it.

OpenStudy (agent0smith):

\[\large \frac{ x }{ \sqrt 5 } = \csc t \]

OpenStudy (loser66):